OpenStudy (anonymous):

The difference between the sum of two numbers and the difference of the two numbers is 6. What is the larger of the two numbers if their product is 15?

OpenStudy (bahrom7893):

(x+y)-(x-y)=6 x*y=15

OpenStudy (anonymous):

Let x = one number Let y = the other number (x + y) - (x - y) = 6 xy = 15 Solve system of two equations with two unknowns.

OpenStudy (anonymous):

How do you solve it?

OpenStudy (anonymous):

Simplify the first equation. x + y - x + y = 6 So what does y equal?

OpenStudy (anonymous):

3 ?

OpenStudy (anonymous):

Yes. Now use the other equation to solve for x.

OpenStudy (anonymous):

xy = 15 x(3) = 15 So x = ?

OpenStudy (anonymous):

Oh i get it thanx